OpenStudy (anonymous):

Solve for X : (9/x) = (3/16)

OpenStudy (anonymous):

Cross Multiply first..

OpenStudy (anonymous):

Oh I didn't think of that! D:

OpenStudy (anonymous):

3x=16*9 x=16*9/3

OpenStudy (anonymous):